In 10 mL of blood sample pH was 7,56. A high amount of acid was added into the blood. After that volume of CO2 in a sample was 6,13mL. pKa is 6,1. Find concentrations of CO2 and HCO3- in blood.

Please, help me, I almost understand how everything should be calculated, but somewhere I keep thinking faulty I guess because can't get the possible answers anyhow...
1
Expert's answer
2014-05-20T06:54:16-0400
Find the amount of substance of CO2. Assume that the volume was measured at ambient conditions (298.15 K, 100 kPa).
n(CO2) = pV/RT = 2.473*10-4 mol
Assume that the only substance responsible for blood pH is CO2. Carbon dioxide dissociates in water according to the following equation:
CO2 + H2O --> H+ + HCO3-
We have a buffer system, and its pH can be calculated from the following equation:
pH = pKa + lg [A-]/[HA]
where [A-] is the concentration of bicarbonate ion HCO3-, and [HA] is the concentration of free CO2.
[A-]/[HA] = 10^(pH - pKa) = 28.84
The total amount of CO2 and HCO3- in the solution is equal to the amount of evolved CO2. Therefore we can find the total concentration of these particles:
c(CO2 + HCO3-) = n(CO2)/V(solution) = 2.473*10-4/0.01 = 0.02473 mol/L
Let x denote the concentration of free CO2 in the solution. Hence the concentration of HCO3- ions is 28.84x
x + 28.84x = 0.02473
x = 8.29*10-4 mol/L
28.84x = 2.39*10-2 mol/L

Answer:
c(CO2) = 8.29*10-4 mol/L
c(HCO3-) = 2.39*10-2 mol/L